use an augmented matrix to solve each system of equations
2y=10
6x+y=-7

Answer by Alan3354(69443)


2y=10
6x+y=-7
----------
0x +2y=10
6x + y=-7
===================
Do it like this one.
---
use an augmented matrix to solve each system of equations
-x+y=9
-4x+y=6
-------------
|-1 +1 -9|
|-4 +1 -6|
-----------------
Det = -1 - (-4) = 3
-----
x*Det = -6 - (-9) = 3
x = 3/Det
x = 1
-----------
-y/Det = 6 - 36 = -30
y = 10